OpenStudy (anonymous):

The solution for a system of two linear equations is (0, 0). Which equation below could be one of the equations in the system? A. B. y = –3x C. y = 4x – 5 D. y = 5 – x

myininaya (myininaya):

find out which equation satisfies (0,0) Pretend I wanted to know if (0,0) was on y=5x+2 all i have to do is replace x with 0 and y with 0 since the point is (0,0)=(x=0,y=0). So like this y=5x+2 0=5(0)+2 0=0+2 0=2 but it is impossible for 0 to equal 2 so (0,0) is not on y=5x+2